Can Some one solve for me this 5 problem.

with Question numbers.

this all Problem. thank you  

13) What force is needed to accelerate a 30.0 lb block at 2.2 m/s? 14) What force is needed to stop a 1500 kg car in 40 m if it is initially moving at 35 miles per hour? 15) A particle is moving in a straight line with constant acceleration a. Derive a general equation for velocity v at position x in terms of initial velocity vo, initial position xo, and acceleration a. 16) You are in the left lane (speeding) at 40 m/s when you notice a police car ahead of you traveling at 30 m/s (the speed limit). When you are 45 m apart you start decelerating at 1.0 m/s2. Will you pass the police car and get a ticket? 17) A car is traveling at 25.0 m/s approaching a bicyclist moving at 7.0 m/s in the same direction. There is 75 m between them when the driver hits the brakes. What is the deceleration needed to avoid a collision?
Answer #1

13) -Masse aalb 12-6078 K ma = 1.3.6078 x 212 1S64Force = mass × acceleration

S = ut + 0.5at^2

V = u + at

V^2 = u^2 + 2aS